Use a reference angle to find the exact value of the expression?
cos (8pi/3)

Answer by lwsshak3(11628)

Use a reference angle to find the exact value of the expression?
cos (8pi/3)
8π/3 is one revolution (6π/3)+(2π/3) which gives an angle of 2π/3 in quadrant II.
Reference angle=π/3 in quadrant II where cos<0.
cos (8pi/3)=cos (π/3)=-1/2